How many 6 digit numbers are there with 3 different digits each of them repeated, if the only digits permitted are members of the set {1,2,3,5,7,8,9,}.
Answer by richard1234(7193)

Pick the 3 digits first. There are 7C3 = 7*6*5/3! = 35 ways to do so.

Given the 3 digits, there are 6!/(2!*2!*2!) = 90 ways to order them to produce different 6-digit numbers.

# of ways = 35*90 = 3150